OpenStudy (anonymous):

help plz i need the answer as soon as posible. Samuel hit the baseball from 2 feet off the grount at an initial vertical velocity of 30 feet per second. How long will it take the ball to hit the ground? A. Write the formula. B. Solve the problem. C. State your answer. Round your answer to the nearest hundredth of a second.

OpenStudy (dan815):

what is acceleration in feet/s^2?

OpenStudy (anonymous):

idk

OpenStudy (dan815):

the outside world uses different measurement system :)

OpenStudy (dan815):

look up your gravity acceleration and tell me

OpenStudy (anonymous):

Formula: h(t) = -16^2 + v0t + h0 (the zero is supposed to be down though, forget what that's called, I think an underpower or something) -16t^2 + 30t + 2 = 0 This will not factor easily so I am going to use QF a = -16 b = 30 c = 2 t = [ -b ±√((b^2 - 4ac))] / (2a) t = [ -30 ±√((30^2 - (4*-16 * 2))] / -32 t = [ -30 +√( 900 + 128) ] / -32 = - .06445...toss this out (no time travel allowed) t = (- 30 - √1028) / -32 = 1.93945 It will take approximately 1.94 seconds for the baseball to hit the ground

OpenStudy (dan815):

so your acceleration is -32 ft/sec^2?

OpenStudy (dan815):